# The Faiss Library

### v3

@@ Abstract
Vector databases typically manage large collections of embedding vectors. Currently, AI applications are growing rapidly, consequently, the number of embeddings that need to be stored and indexed is increasing. The Faiss library is dedicated to vector similarity search, a core functionality of vector databases. Faiss is a toolkit of indexing methods and related primitives used to search, cluster, compress and transform vectors. This paper describes the trade-offs in vector search and the design principles of Faiss in terms of structure, approach to optimization and interfacing. We benchmark key features of the library and discuss a few selected applications to highlight its broad applicability.
@@ 1 Introduction
Since its open-source release in 2017, Faiss has emerged as one of the most popular vector search libraries, boasting 30k GitHub stars and more than 4000 citations of its GPU implementation paper [47]. The Faiss packages have been downloaded 3M times. Major vector database companies, such as Zilliz and Pinecone, either rely on Faiss as their core engine or have reimplemented Faiss algorithms.
@@ 2 Related work
Since the work by [44], quantization-based ANN has emerged as a powerful alternative to binary codes [89]. We refer the reader to the survey by [58] that discusses numerous research works related to quantization-based compact codes.
LSH often refers to indexing with multiple partitions, such as E2LSH [22]. However, we do not consider this scenario as the performance is generally inferior to that of learnt partitions [66]. Early data-aware methods that proved successful on large datasets include multiple partitions based on kd-tree or hierarchical k-means [62]. They are often combined with compressed-domain representation and are especially appropriate for very large-scale settings [43, 44].
@@
Most of the research works on vector search have been open-sourced, and some of these evolved in relatively comprehensive software packages for vector search. For instance, FLANN includes several index types and a distributed implementation described extensively in [62]. The first implementation of product quantization relied on the Yael library [27], that already had a few of the Faiss principles: optimized primitives for clustering methods (GMM and k-means), scripting language interface (Matlab and Python) and benchmarking operators. NMSlib, a package originally designed for text retrieval, was the first to include HNSW [11]. It also provides several index types. The HNSWlib library later became the reference implementation of HNSW [55]. Google’s SCANN library is a thoroughly optimized implementation of IVFPQ [44] on SIMD and includes several index variants for various database scales. SCANN was open-sourced together with [36], which notably omits the engineering optimizations that underpin the library’s remarkable speed. Diskann [79] is Microsoft’s foundational graph-based vector search library, originally built to leverage hybrid RAM/flash memory. It also offers a RAM-only version. It was later extended to perform efficient updates [78], out-of-distribution search [42] and filtered search [34].
Faiss was open-sourced concurrently with the publication [47], which details the GPU implementation of several index types. The present paper complements this previous work by describing the library as a whole.
@@
Techniques based on vector compression, either with binary sketches or quantization, often correspond to estimators of the target distance. For some of these methods, the variance of these estimators can be computed in closed form [16, 44], characterizing the distance estimation error. Again, the quality of the estimator is only a proxy for the actual k-nearest-neighbor problem.
@@ 3 Performance axes of a vector search library
In the equations above, we leave the definition of the distance undefined. The most commonly used distances in Faiss are the L2 distance, the cosine similarity and the inner product similarity (for the latter two, the $\mathrm{argmin}$ should be replaced with an $\mathrm{argmax}$). These measures have useful analytical properties: for example, they are invariant under $d$-dimensional rotations.
The Euclidean distance is the default for comparing vectors [53]. When vectors are obtained by metric learning, the cosine similarity is often used to avoid collapsing or exploding embeddings [23]. Maximum inner product search (MIPS) is most often used for recommendation systems that require to compare user and item embeddings [65].
@@
Table 2: One wants to compare a query vector $x$ and a database vector $y$ given a particular metric (rows). The table indicates how to preprocess $(x,y)\longmapsto(x^{\prime},y^{\prime})$ so that an index with another metric (columns) returns the nearest neighbors for the source metric. Some cases require adding 1 or 2 extra dimensions to the original vectors, as is denoted by the vector concatenation symbol [.;.][.;.][ . ; . ]. The positive scalar parameters $\alpha$ and $\beta$ are arbitrary and chosen to avoid negative values under a square root.
@@ 3.2 Metrics for Approximate Nearest Neighbor Search
Accuracy of k-nearest neighbor search is generally evaluated with “$n$-recall@$k$”, which is the fraction of the $n$ ground-truth nearest neighbors that are in the $k$ first search results (with $n\leq k$). Most often $k$$=$$1$ or $n$$=$$k$ (in which case the measure is a.k.a. “intersection measure”). When $n$ = $k$ = $1$, the recall measure and intersection are the same, and the recall is called “accuracy”. In some publications [44], recall@$k$ means 1-recall@$k$, while in others [77] it corresponds to $k$-recall@$k$.
For range search, the exact search result is obtained by applying (3), using threshold $\varepsilon$. To yield result list $\widehat{R}$, the approximate search uses a (possibly different) threshold $\varepsilon^{\prime}$. Thus, standard retrieval metrics can be computed: precision $P=|R\cap\widehat{R}|/|\widehat{R}|$ and recall $R=|R\cap\widehat{R}|/|R|$. By sweeping $\varepsilon^{\prime}$ from small to large, the result list $\widehat{R}$ increases, producing a precision-recall curve. The area under the PR-curve is the mean average precision score of range search [76]. Setting $\varepsilon^{\prime}\neq\varepsilon$ is relevant when approximate vector representations distort the distance metric (more on this in Section 4).
@@ 3.4 Exploring search-time settings
$\pi^{\prime}\geq\pi\Rightarrow\left\{\begin{array}[]{c}S(\pi^{\prime})\leq S(\pi),\\ A(\pi^{\prime})\geq A(\pi).\\ \end{array}\right.$ (6)
@@
In practice, we evaluate settings from $\mathcal{P}$ in a random order. The pruning becomes more and more effective throughout the process. It is also more effective when the number of parameters is larger. Figure 1 shows an example with $|\mathcal{P}|=5808$ combined parameter settings. The pruning from (9) reduces this to 398 experiments, out of which $|\mathcal{P}^{*}|=87$ are optimal. The Faiss OperatingPoints object implements this.
@@ 4.3 Faiss additive quantization options
  • $\left\|\sum_{i=1}^{m-1}T_{i}[c_{i}]-x\right\|^{2}$ is the encoding error of the previous step $m-1$;
  • $-2\langle T_{m}[j],x\rangle$ is computed on entry to the encoding (it is the only computation complexity that depends on $d$);
@@
Figure 3: Accuracy vs. code size trade-off for different codecs on the Deep1M and Contriever1M datasets. We show Pareto-optimal variants with larger dots and indicate the quantizer in text for some of them. Note that contriever vectors can be encoded to MSE=$2\cdot 10^{-4}$ in 768 bytes with SQ8 (that setting is widely out-of-range for the plot).
@@
This consists in computing distances without decompressing the stored vectors. It is acceptable to perform pre-computations on the query vector $q$ because it is assumed that the cost of these pre-computations will be amortized over many query-to-code distance comparisons.
@@ 4.4 Vector compression benchmark
Figure 3 shows the trade-off between code size and accuracy for many variants of the codecs. Additive quantizers are the best options for small code sizes. For larger code sizes, it is beneficial to independently encode several sub-vectors with product-additive quantizers. LSQ is more accurate than RQ for small codes, but does not scale well to longer codes. Note that product quantizers are a bit less accurate than the additive quantizers but given their low encoding time they remains an attractive option. The scalar quantizers perform well for very long codes and are even faster. The 2-level PQ options are what an IVFPQ index uses as encoding: a first-level coarse quantizer and a second level refinement of the residual (more about this in Section 5.1).
@@ 5 Non-exhaustive search
Non-exhaustive search is the cornerstone of fast search implementations for datasets larger than around $N$=10k vectors. In that case, the aim of the indexing method is to quickly focus on a subset of database vectors that are most likely to contain the search results.
@@ 5.1 Inverted files
Figure 4 shows the optimal settings of ${K_{\mathrm{IVF}}}$ for various database sizes. For a small ${K_{\mathrm{IVF}}}=4096$, the coarse quantization runtime is negligible and the search time increases linearly with the database size. On larger datasets, it is beneficial to increase ${K_{\mathrm{IVF}}}$. As in (18), the ratio ${K_{\mathrm{IVF}}}/\sqrt{N}$ is roughly 15 to 20. Note that this ratio depends on the data distribution and the target accuracy. Interestingly, in a regime where ${K_{\mathrm{IVF}}}$ is larger than the optimal setting for $N$ (e.g. ${K_{\mathrm{IVF}}}=2^{18}$ and $N=$5M), the ${P_{\mathrm{IVF}}}$ needed to reach the target accuracy decreases with the dataset size, and so does the search time. This is because when ${K_{\mathrm{IVF}}}$ is fixed and $N$ increases, for a given query vector, the nearest database vector is either the same or a new one that is closer, so it is more likely to be found in a quantization cluster nearer to the query.
@@
Thus, with this model, the search time increases faster for higher accuracy targets, but $\alpha<0.5$, so the runtime dependence on the database size is below $\sqrt{N}$.
Figure 4: Search time as a function of the database size $N$ for BigANN1B with different ${K_{\mathrm{IVF}}}$ settings. The ${P_{\mathrm{IVF}}}$ is set so that the 1-recall@1 is 90%. The full lines indicate that the coarse quantizer is exact, the dashed lines rely on a HNSW coarse quantizer. For some setting we indicate the ratio ${K_{\mathrm{IVF}}}/\sqrt{N}$
@@
Figure 5: Comparing IVF indexes with and without residual encoding for ${K_{\mathrm{IVF}}}\in\{2^{10},2^{14}\}$ on the Deep1M dataset ($d$=96 dimensions), with different product quantization settings. We measure the recall that can be achieved within 3000 distance comparisons.
@@ 5.3 How to choose an index
Figure 8: Comparison between the Faiss and SCANN libraries indexing Deep10M in the ann-benchmarks setup (batch mode).
@@ 5.5 Comparison with other libraries
ANN-benchmarks [3] is a codebase that compares several ANNS implementations. We use this setup to compare Faiss with SCANN [36], another industry-standard package for vector search. We run the search in the following setting: batch search on the Deep10M dataset with 10-recall@10 as the metric, where the training is performed on the database vectors. This differs slightly from the evaluation protocol used originally for this dataset.
Figure 8 shows that Faiss is 1.5$\times$ to 4$\times$ faster than SCANN, depending on the operating point. For Faiss we used IVF with a Product Residual Quantizer optimized for SIMD (see Appendix A.3), followed by re-ranking (Section 3.5).
@@ 6 Database operations
In the experiments above, the indexes are built in one go with all the vectors, while search operations are performed with one batch containing all query vectors. In real settings, the index evolves over time, vectors may be dynamically added or removed, searches may have to take into account metadata, etc. In this section we show how Faiss supports some of these operations, mainly on IVF indexes. Specific APIs are available to interface with external storage (Appendix A.4) if fine-grained control is required.
@@ 6.1 Identifier-based operations
Faiss indexes support two types of identifiers: sequential and arbitrary ids. Sequential ids are based on the order of additions in the index. Alternatively, the user can provide arbitrary 63-bit integer ids associated to each vector. The corresponding addition methods for the index are add and add_with_ids. In addition, Faiss supports removing vectors (remove_ids) and updating them (update_vectors) by passing the corresponding ids.
Unlike e.g. Usearch [87], Faiss does not store arbitrary metadata with the vectors, only 63-bit integer ids can be used (the sign bit is reserved for invalid results).
@@
Sequential indexes (IndexFlatCodes) store vectors as a flat array. They support only sequential ids. When arbitrary ids are needed, the index can be embedded in a IndexIDMap, that translates sequence numbers to arbitrary ids using a int64 array. This enables add_with_ids and returns the arbitrary ids at search time. For more id-based operations, a more complete wrapper, IndexIDMap2, maps arbitrary ids back to sequential ids using a hash table.
As graph indexes rely on an embedded IndexFlatCodes to store the actual vectors, they should also be wrapped with the mapping classes. However, HNSW does not support suppression and mutation, and NSG does not even support adding vectors incrementally. Supporting this requires heuristics to rebuild the graph when it is mutated, which are implemented in HNSWlib and [78] but are suboptimal indexing-wise.
@@
The IVF indexing structure supports user-provided ids natively at addition and search time. However, id-based access may require a sequential scan, as the entries are stored in an arbitrary order in the inverted lists. Therefore, the IVF index can optionally maintain a DirectMap, which maps user-visible ids to the inverted list and the offset they are stored in. It supports lookup, removal and update by ids. The map can be an array, which is appropriate for sequential ids, or a hash table, for arbitrary 63-bit ids. Obviously, the direct map incurs a memory overhead and an add-time computation overhead, therefore, it is disabled by default.
@@ 6.2 Filtered search
To speed up the predicate, we can use bit manipulations. In this example, $N=10^{7}$, so we use only $\lceil\log_{2}N\rceil=24$ bits of the ids, leaving $63-24=39$ bits that are always 0. We associate to each word $j$ a 39-bit signature $S[j]$, and to each set of words the binary “or” of these signatures. The query is represented by $s_{\mathrm{q}}=S[w_{1}]\vee S[w_{2}]$. Database entry $i$ with words $W_{i}$ is represented by $s_{i}=\vee_{w\in W_{i}}S[w]$. Then the following implication holds: if $\{w_{1},w_{2}\}\subset W_{i}$ then all 1 bits of $s_{\mathrm{q}}$ are also set to 1 in $s_{i}$:

### v4

@@ Abstract
Vector databases typically manage large collections of embedding vectors. As AI applications are growing rapidly, the number of embeddings that need to be stored and indexed is increasing. The Faiss library is dedicated to vector similarity search, a core functionality of vector databases. Faiss is a toolkit of indexing methods and related primitives used to search, cluster, compress and transform vectors. This paper describes the trade-offs in vector search and the design principles of Faiss in terms of structure, approach to optimization and interfacing. We benchmark key features of the library and discuss a few selected use cases to highlight its broad applicability.
@@ 1 Introduction
Since its open-source release in 2017, Faiss has emerged as one of the most popular vector search libraries, boasting 37k GitHub stars and more than 5200 citations of its GPU implementation paper [47]. The Faiss packages have been downloaded 6M times. Major vector database companies, such as Zilliz and Pinecone, either rely on Faiss as their core engine or have reimplemented Faiss algorithms.
@@ 2 Related work
Since the original work introducing product quantization search [44], quantization-based ANN has emerged as a powerful alternative to binary codes [90]. We refer the reader to the survey by [58] that discusses numerous research works related to quantization-based compact codes.
The term LSH can also refer to indexing with multiple partitions, such as E2LSH [22]. However, we do not consider this scenario as the performance is generally inferior to that of learnt partitions [66]. Early data-aware methods that proved successful on large datasets include multiple partitions based on kd-tree or hierarchical k-means [62]. They are often combined with compressed-domain representation and are especially appropriate for very large-scale settings [43, 44].
@@
Most of the research works on vector search have been open-sourced, and some of these evolved in relatively comprehensive software packages for vector search. For instance, FLANN includes several index types and a distributed implementation described extensively in [62]. The first implementation of product quantization relied on the Yael library [27], that already had a few of the Faiss principles: optimized primitives for clustering methods (GMM and k-means), scripting language interface (Matlab and Python) and benchmarking operators. NMSlib, a package originally designed for text retrieval, was the first to include HNSW [11]. It also provides several index types. The HNSWlib library later became the reference implementation of HNSW [55]. Google’s SCANN library is a thoroughly optimized implementation of IVFPQ [44] on SIMD and includes several index variants for various database scales. SCANN was open-sourced but the accompanying paper [36] omits to mention the engineering optimizations that underpin the library’s remarkable speed. DiskANN [79] is Microsoft’s foundational graph-based vector search library. It was originally built to leverage hybrid RAM/flash memory, but now offers a RAM-only version and was later extended to perform efficient updates [78], out-of-distribution search [42] and filtered search [34].
Faiss was open-sourced concurrently with a publication [47] that details the GPU implementation of several index types. The present paper complements this previous work by describing the library as a whole.
@@
Techniques based on vector compression, either with binary sketches or quantization, often correspond to estimators of the target distance. For some of these methods, the variance of the estimator can be computed in closed form [16, 44], characterizing the distance estimation error. Again, the quality of the estimator is only a proxy for the actual k-nearest-neighbor problem.
@@ 3 Performance axes of a vector search library
In the equations above, we leave the definition of the distance undefined. Faiss aims at supporting a range of distance metrics, but in order to avoid overloading the library, it leaves out metrics that are equivalent up to a monotonous transformation of distances (like Euclidean vs. squared Euclidean) or up to a preprocessing of the input vectors (like Mahalanobis or cosine, see below). The full list of metrics is in Appendix A.6.
The most commonly used distances in Faiss are the L2 distance, the cosine similarity and the inner product similarity (for the latter two, the $\mathrm{argmin}$ should be replaced with an $\mathrm{argmax}$). These measures have useful analytical properties: for example, they are invariant under $d$-dimensional rotations.
In terms of applications, the Euclidean distance is the default for comparing vectors [53]. When vectors are obtained by metric learning, the cosine similarity is often used to avoid collapsing or exploding embeddings [23]. Maximum inner product search (MIPS) is most often used for recommendation systems that require to compare user and item embeddings [65].
@@
Table 2: One wants to compare a query vector $x$ and a database vector $y$ given a particular metric (rows). The table indicates how to preprocess $(x,y)\longmapsto(x^{\prime},y^{\prime})$ so that an index with another metric (columns) returns the nearest neighbors for the source metric. Some cases require adding 1 or 2 extra dimensions to the original vectors, as is denoted by the vector concatenation symbol $[.;.]$. The positive scalar parameters $\alpha$ and $\beta$ are arbitrary and chosen to avoid negative values under a square root.
@@ 3.2 Metrics for Approximate Nearest Neighbor Search
Accuracy of k-nearest neighbor search is generally evaluated with “$n$-recall@kk”, which is the fraction of the $n$ ground-truth nearest neighbors that are in the $k$ first search results (with $n\leq k$). Most often kk==11 or nn==$k$ (in which case the measure is a.k.a. “intersection measure”). When $n$ = $k$ = $1$, the recall measure and intersection are the same, and the recall is called “accuracy”. In some publications [44], recall@$k$ means 1-recall@$k$, while in others [77] it corresponds to $k$-recall@$k$.
[…]
@@ 3.4 Exploring search-time settings
[…]
@@
In practice, we evaluate settings from $\mathcal{P}$ in a random order. The pruning becomes more and more effective throughout the process. It is also more effective when the number of parameters is larger. Figure 1 shows an example with $|\mathcal{P}|=5808$ combined parameter settings. The pruning from (9) reduces this to 398 experiments, out of which $|\mathcal{P}^{*}|=87$ are optimal. The Faiss OperatingPoints object implements this pruning.
@@ 4.3 Faiss additive quantization options
  • $\left\|\sum_{i=1}^{m-1}T_{i}[c_{i}]-x\right\|^{2}$ is the encoding error of the previous step m1m-1;
  • $2\langle T_{m}[j],x\rangle$ is computed on entry to the encoding (it is the only computation complexity that depends on $d$);
@@
[…]
@@
The distances are computed without decompressing the stored vectors. It is acceptable to perform pre-computations on the query vector $q$ because the cost of these pre-computations is amortized over many query-to-code distance comparisons.
@@ 4.4 Vector compression benchmark
Figure 3 shows the trade-off between code size and accuracy for many variants of the codecs. Additive quantizers are the best options for small code sizes. For larger code sizes, it is beneficial to independently encode several sub-vectors with product-additive quantizers. LSQ is more accurate than RQ for small codes, but does not scale well to longer codes. Note that product quantizers are a bit less accurate than additive quantizers, but given their low encoding time they remain an attractive option. The scalar quantizers perform well for very long codes and are even faster. The 2-level PQ options are what an IVFPQ index uses as encoding: a first-level coarse quantizer and a second level refinement of the residual (more about this in Section 5.1).
@@ 5 Non-exhaustive search
Non-exhaustive search is the cornerstone of fast search implementations for datasets larger than around NN=10k vectors. In that case, the aim of the indexing method is to quickly focus on a subset of database vectors that are most likely to contain the search results.
@@ 5.1 Inverted files
Figure 4 shows the optimal settings of $K_{\mathrm{IVF}}$} for various database sizes. For a small ${K_{\mathrm{IVF}}}=4096$, the coarse quantization runtime is negligible and the search time increases linearly with the database size. On larger datasets, it is beneficial to increase $K_{\mathrm{IVF}}$}. As in (18), the ratio ${K_{\mathrm{IVF}}}/\sqrt{N}$ is roughly 15 to 20. Note that this ratio depends on the data distribution and the target accuracy. Interestingly, in a regime where $K_{\mathrm{IVF}}$} is larger than the optimal setting for $N$ (e.g. ${K_{\mathrm{IVF}}}=2^{18}$ and N=N=5M), the $P_{\mathrm{IVF}}$} needed to reach the target accuracy decreases with the dataset size, and so does the search time. This is because when $K_{\mathrm{IVF}}$} is fixed and $N$ increases, for a given query vector, the nearest database vector is either the same or a new one that is closer, so it is more likely to be found in a quantization cluster nearer to the query.
@@
Thus, with this model, the search time increases faster for higher accuracy targets, but $\alpha<0.5$, so the runtime dependence on the database size is below $\sqrt{N}$. This empirical complexity analysis has been used before for graph-based indices [92].
Figure 4: Search time as a function of the database size NN for BigANN1B with different $K_{\mathrm{IVF}}$} settings. The $P_{\mathrm{IVF}}$} is set so that the 1-recall@1 is 90%. The full lines indicate that the coarse quantizer is exact, the dashed lines rely on a HNSW coarse quantizer. For some setting we indicate the ratio ${K_{\mathrm{IVF}}}/\sqrt{N}$
@@
Figure 5: Comparing IVF indexes with and without residual encoding for ${K_{\mathrm{IVF}}}\in\{2^{10},2^{14}\}$ on the Deep1M dataset (dd=96 dimensions), with different product quantization settings. We measure the recall that can be achieved within 3000 distance comparisons.
@@ 5.3 How to choose an index
Figure 8: Comparison between the Faiss two other vector search libraries (SCANN and Diskann/vamana) when Deep10M in the ann-benchmarks setup (batch mode).
@@ 5.5 Comparison with other libraries
ANN-benchmarks [3] is a codebase that compares several ANNS implementations. We use this setup to compare Faiss with SCANN [36] and Vamana [79], two other industry-standard packages for vector search. We run the search in the following setting: batch search on the Deep10M dataset with 10-recall@10 as the metric, where the training is performed on the database vectors. This differs slightly from the evaluation protocol used originally for this dataset.
Figure 8 shows that Faiss is faster than SCANN and about on-par with Vamana, depending on the operating point. For Faiss we used IVF with a Product Residual Quantizer optimized for SIMD (see Appendix A.3), followed by re-ranking (Section 3.5).
@@ 6 Database operations
In the experiments above, the indexes are built in one go with all the vectors, while search operations are performed with one batch containing all query vectors. In real settings, the index evolves over time, vectors may be dynamically added or removed, searches may have to take into account metadata, etc. In this section we show how Faiss supports some of these operations. Specific APIs are available to interface with external storage (Appendix A.4) if fine-grained control is required.
@@ 6.1 Identifier-based operations
Faiss indexes support two types of identifiers: sequential and arbitrary ids. Sequential ids are based on the order of additions in the index. Alternatively, the user can provide arbitrary 63-bit integer ids associated to each vector (the sign bit is reserved for invalid results). The corresponding addition methods for the index are add and add_with_ids. Unlike e.g. Usearch [87], Faiss does not store arbitrary metadata with the vectors.
Index updates.
The Faiss API includes methods to remove vectors (remove_ids) and update them (update_vectors) by passing the corresponding ids.
Note that if the vector distribution changes significantly because of additions/removals or updates, then the efficiency of any technique that fits the data distribution degrades. This includes IVF and PQ compression. This degradation can be addressed by explicit updates of the index structure [8].
@@
Sequential indexes (IndexFlatCodes) store vectors as an array. They support only sequential ids. When arbitrary ids are needed, the index can be embedded in a IndexIDMap, that translates sequence numbers to arbitrary ids using a int64 array. This enables add_with_ids and returns the arbitrary ids at search time. IndexIDMap2 in addition maps arbitrary ids back to sequential ids using a hash table. The graph indexes rely on an embedded IndexFlatCodes to store the actual vectors. To use them with non-sequential ids, they should be wrapped with the same mapping objects.
@@
The IVF indexing structure supports user-provided ids natively at addition and search time. However, id-based access may require a sequential scan, as the entries are stored in an arbitrary order in the inverted lists. Therefore, the IVF index can optionally maintain a DirectMap, which maps user-visible ids to the inverted list and the offset they are stored in. It supports lookup, removal and update by ids. The map can be an array, which is appropriate for sequential ids, or a hash table, for arbitrary 63-bit ids. The direct map incurs a memory overhead and an add-time computation overhead, therefore, it is disabled by default.
Graph indexes.
The graph index HNSW does not support suppression and mutation, and NSG does not even support adding vectors incrementally. Supporting this requires heuristics to rebuild the graph when it is mutated, which are implemented in HNSWlib and FreshDiskANN [78] but suboptimal indexing-wise.
@@ 6.2 Filtered search
To speed up the predicate, we can use bit manipulations. In this example, $N=10^{7}$, so we use only $\lceil\log_{2}N\rceil=24$ bits of the ids, leaving 6324=3963-24=39 bits that are always 0. We associate to each word $j$ a 39-bit signature $S[j]$, and to each set of words the binary “or” of these signatures. The query is represented by $s_{\mathrm{q}}=S[w_{1}]\vee S[w_{2}]$. Database entry $i$ with words $W_{i}$ is represented by $s_{i}=\vee_{w\in W_{i}}S[w]$. Then the following implication holds: if $\{w_{1},w_{2}\}\subset W_{i}$ then all 1 bits of $s_{\mathrm{q}}$ are also set to 1 in $s_{i}$: